### v1

# Numaan NaeemKaushal Kumar MauryaKseniia PetukhovaEkaterina Kochmar

Abstract
We present AITutor-EvalKit, an application that uses language technology to evaluate the pedagogical quality of AI tutors, provides software for demonstration and evaluation, as well as model inspection and data visualization. This tool is aimed at education stakeholders as well as *ACL community at large, as it supports learning and can also be used to collect user feedback and annotation.
[Uncaptioned image]AITutor-EvalKit: Exploring the Capabilities of AI Tutors
Numaan Naeemthanks: Equal contribution.   Kaushal Kumar Maurya   Kseniia Petukhova   Ekaterina Kochmar Mohamed bin Zayed University of Artificial Intelligence (MBZUAI) {Numaan.Naeem,Kaushal.Maurya,Kseniia.Petukhova,Ekaterina.Kochmar}@mbzuai.ac.ae
1 Introduction
Personalized one-on-one tutoring has long been recognized as a highly effective educational approach Bloom (1984). Yet, its widespread adoption is constrained by the limited availability of qualified tutors Wang et al. (2024b) and the high costs associated with tutor training Kelly et al. (2020), among other impediments Yoon et al. (2007); Boyd et al. (2008). An alternative to human tutoring is provided by AI tutoring systems, especially those relying on recent advances in large language models (LLMs), such as Khanmigo Khan (2024) and Tutory. [^1] https://tutory.io Despite the remarkable success of LLMs in various tasks Minaee et al. (2024), their adoption in education is hindered by lack of a clear understanding of what these models are capable of Tack et al. (2023); Jurenka et al. (2024) and how pedagogically useful they are Macina et al. (2023b), which results in lack of trust on the part of key educational stakeholders. With the fast development of LLMs and their easy integration into learning tools, questions about the evaluation of AI-driven tutor performance become increasingly relevant Kosmyna et al. (2025). The goal of our tool is two-fold: (1) via the open-source and open-access code, we provide a practical, customizable and versatile evaluation tool that can be applied to a variety of educational scenarios; and (2) via informative demonstrations, we aim to raise awareness of the stakeholders (e.g., students and teachers) as well as practitioners interested in the state of the art in AI in Education (i.e., *ACL community at large) of the current capabilities of LLMs-as-tutors. Although our tool is an early research prototype, through its extendable nature, we aim to facilitate research in this exciting and emerging area of applied NLP research.
Figure 1: This example shows a sample dialogue and its pedagogical-ability evaluation by the LoMTL model using the AITutor-EvalKit. The evaluation follows the four dimensions from Kochmar et al. (2025): MI (Mistake Identification), ML (Mistake Location), PG (Providing Guidance), and AC (Actionability). TSE: To some extent.
The current version of our tool focuses on the pedagogical quality evaluation of tutor responses in the context of Student Mistake Remediation (SMR) Boaler (2013); Handa et al. (2023) in the mathematical domain, where tutors address errors or misconceptions that hinder student progress Wang et al. (2024a, b); Macina et al. (2023a). As the foundation for evaluation, we use the established taxonomy from Maurya et al. (2025), which is grounded in the learning sciences principles and which allows us to assess the quality of tutor responses along four key SMR dimensions Kochmar et al. (2025): (1) mistake identification, concerned with whether tutor’s response notifies the student of the committed mistake; (2) mistake location, focusing on whether the tutor clearly points to the erroneous part in the student’s solution; (3) providing guidance, evaluating the quality of the pedagogical guidance; and (4) actionability, assessing whether tutor’s response makes it clear what the student should do next. These dimensions are further outlined in Table 2, with an illustrative example provided in Figure 1.
The structure and implementation of the front- and backend of our tool are described in Section 3. Using MRBench dataset Maurya et al. (2025) and taking inspiration from the BEA 2025 shared task on AI tutor response evaluation Kochmar et al. (2025), we introduce a novel, efficient, and light-weight multi-task learning model that addresses the four dimensions of pedagogical quality evaluation ($\S$ 3.1). The outputs of the model, as well as those of an open-source (Prometeus Kim et al. (2024)) and a commercial (GPT-5) LLMs used as judges, are displayed using an interactive browser-based UI with helpful visualizations ($\S$ 3.2). The evaluation results (Section 4) suggest that while our model achieves competitive results when evaluated against gold standard annotation, its outputs are also perceived by users to be at least as accurate as those of the commercial LLM-as-judge models, and the UI is considered informative and easy to use. We present and publicly release:
  • The first of its kind, open-access and open-source model aimed at evaluation of the pedagogical quality of AI tutor responses available at MIT-licensed python repository: https://github.com/kaushal0494/AITutor-EvalKit. We believe it to be useful for AI-in-Education practitioners and developers, as it is highly customizable, allowing researchers to apply it to further educational contexts and dialogues, as well as to extend it to other scenarios and domains.
  • An interactive UI available at https://demo-ai-tutor.vercel.app, which communicates the results and showcases the capabilities of AI tutors in an interpretable way, which we consider to be of interest to education stakeholders and the *ACL community at large. The demo tool can also be run locally with the user’s own data and models.
  • A short video demonstrating the tool available at https://www.youtube.com/watch?v=9qgDfrhzOvg.
2 Related Work
Over the years, the NLP community has seen significant advances in the development of publicly available toolkits for modeling and evaluation, including Hugging Face Transformers Wolf (2019), NLTK Bird (2006), and Scikit-learn Pedregosa et al. (2011). These toolkits have enhanced code reusability, enabling researchers to focus on developing more sophisticated models and metrics. However, in the educational domain, especially in conversational dialogues, there remains a lack of robust tools to push research boundaries. Some progress has been made with frameworks like ConvoKit Chang et al. (2020), which facilitates the analysis of general conversational data. More recently, Edu-ConvoKit Wang and Demszky (2024) was developed for preprocessing, annotation, and analysis of educational dialogues. While these toolkits contribute significantly to data handling and analysis, they fall short of addressing the evaluation of pedagogical quality in AI-driven educational systems. To the best of our knowledge, there is no toolkit that focuses on the on-the-fly assessment of the pedagogical abilities of AI tutors. With AITutor-EvalKit, we aim to fill this critical gap by providing an open-source toolkit that integrates popular taxonomies like that of Maurya et al. (2025) into a single unified framework for the systematic evaluation of AI tutors and can be extended with further approaches such as that of Macina et al. (2025), thereby fostering progress in this underexplored yet important research area.
Figure 2: AITutor-EvalKit pipeline: The backend module includes several model options to assess the pedagogical soundness of tutors’ responses, and the frontend projects evaluation outputs via an interactive user interface.
3 System and Demonstration Description
AITutor-EvalKit consists of two major modules: backend and frontend, with the pipeline illustrated in Figure 2. The backend module includes several models to evaluate tutor responses, and frontend seamlessly integrates these evaluation outputs with an interactive, customizable, and flexible UI. Different audience groups can benefit from different functionalities of the toolkit: e.g., researchers and developers can use both modules, train their own automated models, use their own datasets, and launch the demo UI locally, while teachers and policymakers can use the frontend module to understand the capabilities of LLMs-as-tutors and make decisions accordingly.
3.1 Backend Module: Evaluation Models
The backend module has two components: automated evaluation and evaluation by LLM-as-a-judge. In this section, we provide details on the training data (for the automated evaluation model), test data, and automated evaluation model itself, as well as the usage of open-source LLMs and closed-source GPT models as judges.
3.1.1 Dataset
As discussed in Section 1, we used MRBench dataset by Maurya et al. (2025), which has 491 dialogues (300 in the development set and 191 in the test set), each paired with seven LLM-generated tutor responses and two human tutor responses. Each response is annotated by human annotators for each of the four dimensions using the categories “Yes,” “To some extent,” and “No.” We also created a randomly selected demonstration set consisting of 10 dialogues, which is a subset of the test set. More details on the dimensions, annotation, and data statistics can be found in Appendix A.
3.1.2 Automated Evaluation: LoMTL
This component implements a ternary classifier to evaluate tutor responses across the four considered dimensions. The models developed by participants in the BEA shared task provide a good starting point, as summarized by Kochmar et al. (2025). However, upon closer inspection, we found that several teams relied on closed-source models, fully fine-tuned models, LoRA-tuned models, and some even used multiple models or ensembles for each dimension, which makes these approaches difficult to scale across all dimensions.
Considering this, we propose a novel LoRA-based multi-task model (called LoMTL) that fine-tunes a single small LLM in a LoRA setting across all four dimensions, each treated as a task in a multi-task learning setting. This modeling is naturally suited to the four tasks, all of which are ternary, closely related classification problems, allowing them to benefit from shared learning during training. Additionally, we use sampling and balanced batching methods to improve performance. We observe that a small google/gemma-2-2b-it Team et al. (2024) model using LoMTL achieves performance competitive with the top-performing teams in the shared task while remaining highly efficient and scalable. Appendix B presents more details on model development, prompts, configuration, and comparative results with the BEA shared task’s top-performing teams.
3.1.3 LLM-as-a-Judge Evaluation
This component provides functionality for evaluation using LLMs-as-judges. Following Maurya et al. (2025), we have selected Prometheus2 Kim et al. (2024) as the primary open-source LLM. However, the implementation of this module is flexible enough to support the family of Llama and other open-source causal LLMs for evaluation. Additionally, we have included an option to evaluate using the closed-source OpenAI GPT-5 model, although the implementation supports any model from the GPT family. Users must provide their OpenAI API key to run evaluations.
3.1.4 Flexible Design
Each of the three evaluation setups (automated evaluation, evaluation with open-source LLMs, and evaluation with closed-source GPT models) provides high degree of flexibility in the choice of the base LLM, prompting strategy, and hyper-parameter tuning. For example, each evaluation setup has its own prompting file and configuration, allowing users to customize and use the components of the tool as per their needs. Evaluation can be run using the following short commands:
# Training LoMTL model
cd src/autoeval && ./lora_finetune_runner.sh
# Evaluation with automated (i.e., LoMTL) model
cd src/autoeval && ./lora_evaluation_runner.sh
# Evaluation with open-source LLM
cd src/llmeval &&
python run_open_llm_as_judge_evaluation.py
# Evaluation with GPT5
cd src/llmeval && ./gpt5_eval_runner.sh
3.2 Frontend Module: Demo App
We present an interactive prototype built on top of MRBench, designed to evaluate the pedagogical abilities of AI tutors in educational dialogues. Users can explore the demo in two modes: (1) a Static Mode, where educators can access a deployed version containing 10 conversations from the MRBench test set evaluated by our fine-tuned model and share feedback on the helpfulness of tutor responses; and (2) a Customized Mode, where developers can run the interface locally to analyze their own datasets using either our LoMTL model or their own one, supported by the provided code. The UI consists of three key modules – Automated Evaluation, LLM Evaluation, and Visualizer, that together enable exploration, analysis, and visualization of educational dialogues.
  • Automated Evaluation provides automated evaluation results using our LoMTL model on 4 evaluation dimensions.
  • LLM Evaluation leverages LLMs-as-judges for human-aligned evaluations.
  • Visualizer enables rich visual analytics for interpreting evaluation scores across 4 pedagogical dimensions on the MRBench development set.
The UI supports the selection of evaluation methods and provides instant visualization of performance metrics through plots, bar charts, and spider graphs.
3.2.1 Automated Evaluation UI
The Automated Evaluation module presents results generated by our LoMTL model, assessing AI tutors across four pedagogical dimensions. As shown in Figure 6, it begins with an overview panel summarizing key statistics such as the number of topics, models, dimensions, and conversations. Users can evaluate a single tutor or compare two tutors side by side.
Single Tutor Evaluation: In this mode, users select a problem topic to view the complete student-tutor dialogue in the “Context” block (Figure 7). A drop-down menu allows selection of a tutor model, and the corresponding response is displayed in the “Tutor Response” block. Users can rate the usefulness of the response (Helpful, To Some Extent, or Not Helpful) and optionally view the ground truth solution for better context. Upon clicking “Get Auto-Evaluation Results,” the system generates performance evaluations across the chosen dimensions, with results downloadable in PNG, JPG, or JSON format. A “Best Performance by Dimension” panel highlights the top tutor(s) for each dimension, helping users quickly identify their pedagogical strengths.
Tutor Comparison Mode: This mode allows users to directly compare two tutors by enabling the “Tutor Comparison Mode” option. The selected tutors’ responses are displayed side by side (Figure 8), and users can provide quick feedback indicating which tutor performed better or mark both as equally good or bad. After choosing the evaluation dimensions and clicking “Compare Tutor Responses,” the interface presents a detailed two-column comparison of scores across all selected dimensions. To facilitate interpretation, the “Comparison Visualization” panel (Figure 9) provides four interactive views: the Summary view highlights the leading tutor for each evaluation dimension and identifies the overall winner; the Spider Chart offers a radar-style visualization comparing performance patterns across dimensions; the Bar Chart displays side-by-side scores for each dimension; and the Differences view illustrates the magnitude of score gaps between tutors. All visualizations can be exported in PNG or JPG formats for reporting or analysis. Finally, the “Best Performance by Dimension” panel summarizes the comparative strengths of the tutor pair, providing a concise overview of pedagogical differences. This mode supports structured, interpretable, and visually grounded benchmarking of AI tutor performance.
3.2.2 LLM Evaluation UI
The LLM Evaluation module enables advanced pedagogical assessment of AI tutor responses using LLMs as judges. It extends the functionality of the automated evaluation pipeline by leveraging LLMs to judge and compare tutor responses across four pedagogical dimensions. The UI supports three evaluation modes: single tutor evaluation, tutor model comparison, and judge LLM comparison. As shown in Figure 10, the overview panel summarizes available topics, conversations, tutor models, evaluation dimensions, and judge LLMs, providing a quick snapshot of the evaluation setup. Currently, two LLM judges are supported: GPT-5 and Prometheus-7B-v2.0 Kim et al. (2024).
Single Tutor Evaluation: This mode allows users to analyze how a selected tutor performs on a specific problem using an LLM as a judge. After selecting the problem, tutor, and LLM judge, users can generate dimension-wise evaluation results that reflect the LLM’s assessment of the tutor’s pedagogical performance. A “Best Performance by Dimension” panel highlights the top-performing tutor(s) for each dimension.
Tutor Comparison Mode: This mode enables side-by-side comparison of two tutor models on the same problem, judged by a selected LLM. Tutor responses are displayed together (Figure 11), and upon comparison, the system presents dimension-wise evaluations and visualizations – such as Summary, Spider Chart, Bar Chart, and Differences – to clearly show relative strengths across pedagogical aspects.
Judge Comparison Mode: This mode compares how different LLM judges evaluate the same tutor response. The tutor’s response is displayed once, and evaluations from both judges are presented in parallel with corresponding visualizations. This feature helps assess consistency between LLM judges and identify possible biases in their evaluations.
Together, these modes enable fine-grained, interpretable analysis of tutor behavior, offering insights into both model performance and evaluation reliability across different judging LLMs.
3.2.3 Visualizer UI
The Visualizer module provides a high-level overview of the MRBench development set across four evaluation dimensions through intuitive visual analytics. It begins with a “Dataset Overview” panel summarizing key statistics, including the number of conversations, tutor models, and evaluation dimensions. The module includes three main visualization panels: Tutor Performance Summary, Visualization Controls, and Dataset Visualization.
The Tutor Performance Summary panel presents average scores for each tutor model across all four dimensions, where “Yes,” “To some extent,” and “No” correspond to 1.0, 0.5, and 0.0, respectively (Figure 12). The Visualization Controls panel allows users to select specific tutors and dimensions to generate detailed visualizations (Figure 13). The Dataset Visualization panel then displays the results through spider and bar charts, where spider plots summarize tutor strengths and weaknesses across dimensions (Figure 14), while bar charts show detailed score distributions for selected dimensions (Figure 15), along with averages for each response label.
This visual analytics module enables users to explore and interpret pedagogical quality effectively, supporting comparative analysis and data-driven insights. The same functionality is also available in the customized mode, allowing developers to visualize and analyze their own datasets locally in a similar way.
4 Evaluation
To evaluate the system, we first measure its performance using the metrics adopted from the BEA 2025 shared task Kochmar et al. (2025) – accuracy and macro-F1, and then conduct a human evaluation study, in which participants evaluate both the prediction quality of the LoMTL evaluation model and the usability of our demo tool and UI.
4.1 Intrinsic Evaluation: Quantitative Analysis
For intrinsic evaluation, we compare our LoMTL model’s predictions on the test set from Kochmar et al. (2025) with the gold human annotations and also run GPT-5 and Prometheus2 on the same data. The average results across evaluation dimensions, presented in Table˜1, show that Prometheus2 performs substantially worse than both our model and GPT-5. Our model achieves the highest averaged accuracy and macro-F1 on the full test set, and it also performs competitively on the demonstration subset. While GPT-5 achieves a slightly higher averaged macro-F1 on the demonstration subset, the overall trend indicates that our model provides more reliable evaluations than Prometheus2 and performs on par with, or better than, GPT-5. The per-dimension results are provided in Table˜6 of Appendix C.1.
Model Full Test Set Demonstration Set
Accuracy Macro-F1 Accuracy Macro-F1
LoMTL (ours) 0.72 0.60 0.68 0.55
Prometheus2 0.47 0.41 0.41 0.33
GPT-5 0.66 0.57 0.66 0.58
Table 1: Accuracy and macro-F1 scores (averaged across Mistake Identification (MI), Mistake Location (ML), Providing Guidance (PG), and Actionability (AC)) for our LoMTL model, Prometheus2, and GPT-5 on the full test set from Kochmar et al. (2025) and on the demonstration set. Best results are shown in bold.
4.2 Extrinsic Evaluation: Human Study
Participants were given access to the demo website, detailed guidelines, and an evaluation form. The form instructed them to assess three components: the Automated Evaluation tab, the LLM Evaluation tab, and the Visualizer tab, as detailed in Section 3.2. For the first two tabs, participants were asked to explore at least five different dialogues, review at least two tutor responses per dialogue, and use the feedback field to rate each response as helpful, helpful to some extent, or not helpful. They also examined the model’s evaluation for each response across at least one assessment dimension. Additionally, for each dialogue, participants compared at least one pair of tutor responses in comparison mode and indicated which response they considered a better one. After using Automated Evaluation tab, they reported how frequently they agreed with the model’s judgments on a 1-5 scale, both in single-response and comparison modes. In the LLM Evaluation tab, participants evaluated how often they agreed with GPT-5 and Prometheus2, again in both modes. They were also asked which evaluation model they perceived as more accurate: GPT-5 or the model from the first tab (which corresponds to our LoMTL model), and Prometheus2 or the model from the first tab. Finally, in the Visualizer tab, participants explored visualizations for at least two tutors and rated how informative they found them on a 1-5 scale. For every tab, participants also rated the ease of use on a 1-5 scale. The full questionnaire is provided in Appendix C.
A total of 14 participants took part in the study. Their educational background included individuals pursuing a Master’s degree, those holding a Master’s degree, and those holding a PhD. Eleven participants had teaching experience, and eight had prior experience using an AI tutor.
Most participants perceived the LoMTL evaluation model as more accurate in its judgments than both GPT-5 and Prometheus2. As shown in Figure˜3, the majority of participants agreed with the LoMTL model’s assessments more than half of the time in both single-response and comparison modes. A similar trend was observed for GPT-5: most participants agreed with its judgments more than half of the time, although a larger proportion reported agreement only about half of the time. In contrast, participants tended to agree with Prometheus2 less than half of the time in the single-response mode, but more than half of the time in the comparison mode. Overall, most participants rated the ease of use of all tabs as very easy and found the visualizations very informative.
Figure 3: Participants’ responses indicating how often they agreed with the models’ judgments in single-response and comparison modes.
In total, we collected 95 annotations for single-tutor responses and 115 for pairwise comparisons. The analysis of these annotations is provided in Appendix C.3. We do not present it here because evaluating tutor performance is not the focus of this work. Instead, our goal is to demonstrate how our interactive UI can be used for annotation purposes by collecting data that can later be used to train evaluation models or align language models.
5 Future Work
Future work will extend the toolkit by (1) integrating new user dialogues and LLMs in the frontend module to generate real-time responses and evaluations; (2) expanding the evaluation dimensions; and (3) broadening coverage to additional subjects, grade levels, and languages.
Limitations
We acknowledge that our work has several limitations. Below, we summarize the major ones among them.
Domain and grade level:
In this work, we focus on the mathematical tasks at the middle-school level. This decision is motivated by the availability of the data at this level and in this domain, but we plan to extend our work to other domains and levels as we elaborate in Section 5. Moreover, since users can run our tool on their own data, new domains and levels can already be integrated on the user’s side.
Language:
Similarly, our current work focuses on English only. In the future, we hope to extend it to other languages, as we specify in Section 5, while users of our tool can also apply it to data in other languages with their own evaluation models and LLMs-as-judges on their side.
Educational scenario:
Building on previous work in this domain Maurya et al. (2025); Kochmar et al. (2025), we only address student mistake remediation as an educational scenario. While this is one of the most salient and challenging scenarios in educational dialogues, we recognize this as one of the limitations of our work and plan to address it in the future.
Context length:
Similarly, following up on the previous work, our current evaluation approach is limited to a single turn in the dialogue. We acknowledge this as a limitation, and believe that future work should extend single-turn approaches to multiple-turn or full-dialogue ones.
Taxonomy:
We have built our prototype tool around a well-established evaluation taxonomy of Maurya et al. (2025). While using a single taxonomy is a limitation, our code is open-source and can be extended to incorporate other data, dialogues, taxonomies, and evaluation models on the user’s side.
Models:
Finally, via our demo tool, we only showcase a few LLMs-as-tutors and deploy only two LLMs-as-judges. Since our code is open-source and extendable, more tutor models and LLMs-as-judges can be integrated on the user’s side.
Ethical Considerations
As this work is exploratory, we do not anticipate any significant ethical risks associated with it. Moreover, one of our main goals in this research is to raise awareness of the education stakeholders as well as practitioners interested in the state of the art in AI in Education of the current capabilities of LLMs-as-tutors. Through transparent evaluation of such models, we aim to improve their interpretability, which we hope will help avoid potential future risks associated with a wider adoption of these models in education.
Acknowledgments
We are grateful to the Google Academic Research Award (GARA) 2024 for supporting this research.
Appendix A Pedagogical Dimensions and Data
A.1 The Task
The proposed evaluation toolkit targets the student mistake remediation (SMR) task within educational dialogues between a student and a tutor in the mathematical domain. Formally, let the conversation history be defined as $H=\{(T_{1},S_{1}),(T_{2},S_{2}),\ldots,(T_{t},S_{t})\},$ where $T_{i}$ is the tutor’s $i$-th response and $S_{i}$ is the student’s $i$-th response. Let $S_{k}$ denote the student’s most recent $k$ utterances, where $k\in[1,\ldots,t]$, that contain a mistake or confusion. The tutor’s objective is to generate the next response $T_{t+1}$ that best addresses this mistake or confusion. The backend module of the toolkit focuses on evaluating the pedagogical soundness of the $T_{t+1}$ response.
A.2 Evaluation Dimensions
[…]
Dimension Definition Labels Desiderata
Mistake identification Has the tutor identified a mistake in a student’s response?
(1) Yes
(2) To some extent
(3) No
Yes
Mistake location Does the tutor’s response accurately point to a genuine mistake and its location?
(1) Yes
(2) To some extent
(3) No
Yes
Providing guidance
Does the tutor offer correct and relevant guidance, such as an explanation,
elaboration, hint, examples, and so on?
(1) Yes
(2) To some extent
(3) No
Yes
Actionability Is it clear from the tutor’s feedback what the student should do next?
(1) Yes
(2) To some extent
(3) No
Yes
Table 2: An overview of the evaluation taxonomy, associated definitions, annotation labels, and desired labels from Maurya et al. (2025).
A.3 Dataset Details and Statistics
We used the MRBench dataset to develop the toolkit and the automated evaluation model (i.e., LoMTL). The initial version of MRBench, released by Maurya et al. (2025), contains 191 dialogues. This was extended by Kochmar et al. (2025) which results in 300 dialogues in the development set and 191 dialogues in the test set. In this work, we use the extended version of the MRBench dataset.
The dataset is built on top of two public datasets – MathDial Macina et al. (2023a) and Bridge Wang et al. (2024a) – which provide partial conversational histories from secondary and primary school-level mathematics, respectively, along with human tutor responses as $T_{t+1}$. MathDial includes only one expert tutor response, whereas Bridge includes two responses from expert and novice human tutors. Additionally, each dialogue includes seven $T_{t+1}$ responses generated by seven state-of-the-art LLMs-as-tutors, including GPT-4 Achiam et al. (2023), Gemini Reid et al. (2024), Sonnet Anthropic (2024), Mistral Jiang et al. (2023), Llama-3.1-8B and Llama-3.1-405B Dubey et al. (2024), and Phi3 Abdin et al. (2024).
All $T_{t+1}$ responses (whether from human tutors or LLMs) are annotated by human annotators across four selected dimensions with labels "Yes," "To some extent," and "No," as detailed in Table 2. The proposed LoMTL model is trained on the development set and evaluated on the test set; all results presented in this work are reported on the test set. We further split the development set into a 9:1 ratio for training and validation when developing the LoMTL model. All model checkpoints were selected based on validation performance. Finally, we randomly selected a subset of 10 dialogues from the test set as a demonstration set, which is used in the demo app. These details are summarized in Table 3.
Parameters Value/Details
Number of dialogues (Dev / Test / Total) 300  /  191  /  491
Number of tutor responses (Dev / Test / Total) 2,476  /  1,547  /  4,023
Number of tutors (Total) 9
Number of human tutors 2 (1 expert, 1 novice)
Number of LLM tutors 7
LLM tutor models GPT-4, Gemini, Sonnet, Mistral,
Llama-3.1-8B, Llama-3.1-405B, Phi-3
Source datasets MathDial, Bridge
MathDial Expert human tutor only
Bridge Expert and Novice human tutors
Demonstration set size 10 dialogues (from test set)
Table 3: Key details of the extended MRBench dataset.
Appendix B LoMTL Evaluation Model
B.1 Motivation
Building LoMTL has a two-fold motivation: (1) The current human annotation-based pedagogical ability assessments presented by Maurya et al. (2025) are static in nature. They are not scalable to new LLMs and tutoring systems, which are being developed very frequently nowadays. We need a reliable automated evaluation model that can provide pedagogical assessment on-the-fly for new tutors or responses and help track the progress of AI tutor abilities. (2) The BEA shared task Kochmar et al. (2025) is a good starting point for developing an automated evaluation model. More than 50 international teams participated in the challenge and proposed several novel modeling approaches, including diverse prompting strategies, full instruction tuning, LoRA-based finetuning, supervised finetuning, data augmentation, label balancing, ensembling and so on. However, most teams that participated in all four tracks did not develop a unified approach (with the exception of the MSA team Hikal et al. (2025)) and instead used models with a large number of parameters. This hinders the adaptability of these approaches, as their deployment becomes challenging and costly.
These limitations motivated us to develop LoMTL, a lightweight model with only 2 billion parameters, created by training the google/gemma-2-2b-it model using LoRA in a multi-task learning setting. It achieves comparable performance while being significantly more efficient (see Table 4 for comparison). For instance, the top-performing BJTU team Fan et al. (2025) achieved a macro-F1 score of 0.645 using 288 billion parameters (across all four dimensions). In contrast, the LoMTL model achieved 0.60 with only 2 billion parameters, approximately 0.7% of the BJTU model’s parameter count.
Team/Model Macro-F1 # LLMs # Parameters Parameter Size vs. 2B
BJTU Fan et al. (2025) 0.646 4 (Qwen/Qwen2.5-72B) $4\times 72\text{B}=288\text{B}$ $144\times$ larger
MSA Hikal et al. (2025) 0.643 $5\times 4$ (mistralai/Mistral-7B-v0.1) $20\times 7\text{B}=140\text{B}$ $70\times$ larger
BLCU-ICALL An et al. (2025) 0.632 4 (Qwen/Qwen2.5-7B) $4\times 7\text{B}=28\text{B}$ $14\times$ larger
bea-jh Roh and Bang (2025) 0.625 4 (zai-org/glm-4-9b) $4\times 9\text{B}=36\text{B}$ $18\times$ larger
LoMTL (ours) 0.601 1 (google/gemma-2-2b-it) $2\text{B}$ -
Table 4: Comparison of macro-F1 scores and model parameters between the our automated evaluation model (i.e., LoMTL) and the top-performing teams in the BEA shared task across four dimensions (aka. tasks). Note that, for a fair comparison, we include only the teams that participated in all four tracks of the shared task.
B.2 Training and Inference
In this section, we provide details on the training and evaluation of the LoMTL model. Inspired by the success of LoRA-based modeling from the BEA shared task Kochmar et al. (2025) and by the community Mao et al. (2025), we adapted a LoRA-based fine-tuning approach. Since we have a small amount of training data (approximately 2500 examples for each dimension) and the different tasks are somewhat related, the natural modeling choice is multi-task learning, where each evaluation dimension is formulated as a task. This LoRA-based multi-task fine-tuning approach (called LoMTL) resulted in a compact single model and enabled flexibility in model deployment. Further, we experimented with a small google/gemma-2-2b-it model, which resulted in fast inference. We observed two major issues during training: task imbalance and label imbalance. To mitigate task imbalance, we implemented a balance batching where each batch has an uniform number of examples from each task. For label imbalance, we explored several approaches such as focal loss, label sampling, loss weighting, and sampling methods. We obtained the best performance with oversampling where we randomly sample underrepresented examples in the training dataset. Model training and inference were done with a single 48GB A6000 GPU. The best checkpoints were obtained using validation data (10% of the development dataset).
B.3 Prompts and Configurations
Prompt structure and training/evaluation configurations for the LoMTL model are shown in Figure 4 and Table 5, respectively.
Figure 4: Overview of the prompt components, their associated definitions and details, and the final prompt structure used in LoMTL.
Category Parameter Value
Common Settings
Model MODEL_NAME google/gemma-2-2b-it
Task Dimensions DIMENSIONS Mistake_Identification, Mistake_Location, Providing_Guidance, Actionability
Input Length MAX_LENGTH 1024
Prompt include_label_definitions Enabled
Training-Only Settings
Batching BATCH_SIZE 4
Batching GRAD_ACCUM 1
Training Schedule EPOCHS 3
Training Schedule LEARNING_RATE 1e-4
Training Schedule WEIGHT_DECAY 0.1
Logging LOGGING_STEPS 50
Saving SAVE_STEPS 300
Evaluation Cycle EVAL_STEPS 300
Oversampling OVERSAMPLE_METHOD “random”
Metric for best model METRIC_FOR_BEST “eval_loss"
LoRA LORA_R 8
LoRA LORA_ALPHA 16
LoRA LORA_DROPOUT 0.1
Early Stopping EARLY_PATIENCE 5
Early Stopping EARLY_THRESHOLD 0.0
Evaluation-Only Settings
Generation TEMPERATURE 1.0
Table 5: Summary of training and evaluation configurations along with their corresponding parameter names.
Appendix C Evaluation Experiment Details
C.1 Metrics Across Dimensions
The per-dimension results, presented in Table˜6 show that Prometheus2 performs poorly, while our model notably outperforms GPT-5 on the Mistake Identification and Actionability dimensions. However, it underperforms GPT-5 by 3 and 6 percentage points in terms of macro-F1 for Mistake Location and Providing Guidance, respectively. A similar trend is observed in the ten dialogues used for the demonstration.
Model Full Test Set Demonstration Set
Accuracy Macro-F1 Accuracy Macro-F1
MI ML PG AC MI ML PG AC MI ML PG AC MI ML PG AC
LoMTL (ours) 0.86 0.67 0.63 0.70 0.66 0.55 0.54 0.65 0.76 0.69 0.60 0.68 0.57 0.50 0.47 0.65
Prometheus 0.58 0.53 0.31 0.46 0.48 0.42 0.32 0.43 0.48 0.41 0.42 0.32 0.34 0.30 0.38 0.30
GPT-5 0.67 0.68 0.70 0.58 0.53 0.58 0.61 0.55 0.67 0.64 0.67 0.66 0.53 0.56 0.59 0.64
Table 6: Accuracy and macro-F1 scores of our model, Prometheus, and GPT-5 across Mistake Identification (MI), Mistake Location (ML), Providing Guidance (PG), and Actionability (AC) on the full test set from Kochmar et al. (2025) and on the demonstration set. Best results are shown in bold.
C.2 Full Questionnaire
Below are the informed consent forms that participants were required to read and accept, as well as the full questionnaire.
Figure 5: Informed consent form that participants were required to accept before proceeding with their feedback and annotations.
Background
Item Question / Response Options
What is your highest qualification? Response options: Bachelor’s degree; Master’s degree; PhD degree; Other.
Do you have teaching experience (e.g., lecturing, supervising or mentoring students, TA-ing, or similar)? Response options: Yes; No.
Have you ever used an AI tutor before? Response options: Yes; No.
Automated Evaluation
Item Question / Response Options
Instructions This tab gives you an opportunity to select among 10 short dialogues on relatively simple (no higher than middle-school level) math problems, check students’ misconceptions, and explore various human and AI tutor responses. The quality of these responses is evaluated using a fine-tuned evaluation model.
Steps 1) Explore at least 5 different dialogues (the correct answer is also available to help you spot the student’s mistake).
2) For each dialogue, check at least 2 different tutor responses.
3) Rate each response as Helpful, Not Helpful, or To some extent.
4) Select at least one quality dimension to view the model’s assessment of that response.
5) For each dialogue, use the comparison mode at least once.
How many dialogues have you checked? Free-form response.
How often did you agree with this model’s assessment for a single tutor response? Scale: 1 = almost never, 2 = less than half of the time, 3 = about half of the time, 4 = more than half of the time, 5 = almost always.
Response options: 1, 2, 3, 4, 5.
How often did you agree with this model’s ranking of the tutors in the comparison mode? Same scale as above.
Response options: 1, 2, 3, 4, 5.
On a scale from 1 to 5, how easy was it to use the “Automated Evaluation” tab? Scale: 1 = not easy at all, 5 = very easy.
Response options: 1, 2, 3, 4, 5.
Please feel free to give us any further feedback on this tab. Free-form response.
LLM Evaluation
Item Question / Response Options
Instructions This tab provides access to the same dialogues and tutor responses as in Tab 1, but this time they are evaluated using LLMs as judges. You can choose between GPT-5 and Prometheus.
Steps 1) Explore at least 5 different dialogues (they may be the same ones as before).
2) For each dialogue, check at least 2 different tutor responses.
3) Use each LLM-as-judge at least twice on different dialogues.
4) Select at least one quality dimension for each response.
5) For each dialogue, use the comparison mode at least once.
How many dialogues have you checked? Free-form response.
How often did you agree with GPT-5’s assessment for a single tutor response? Scale: 1 = almost never, 2 = less than half of the time, 3 = about half of the time, 4 = more than half of the time, 5 = almost always.
Response options: 1, 2, 3, 4, 5.
How often did you agree with GPT-5’s ranking of the tutors in the comparison mode? Same scale as above.
Response options: 1, 2, 3, 4, 5.
How often did you agree with Prometheus’ assessment for a single tutor response? Same scale as above.
Response options: 1, 2, 3, 4, 5.
How often did you agree with Prometheus’ ranking of the tutors in the comparison mode? Same scale as above.
Response options: 1, 2, 3, 4, 5.
Which evaluation model did you perceive to be more accurate in its judgments of tutor responses – GPT-5 or the fine-tuned model from Tab 1 (“Automated Evaluation”)? Response options: The model from Tab 1; GPT-5; Hard to say: they perform similarly.
Which evaluation model did you perceive to be more accurate in its judgments of tutor responses – Prometheus or the fine-tuned model from Tab 1 (“Automated Evaluation”)? Response options: The model from Tab 1; Prometheus; Hard to say: they perform similarly.
On a scale from 1 to 5, how easy was it to use the “LLM Evaluation” tab? Scale: 1 = not easy at all, 5 = very easy.
Response options: 1, 2, 3, 4, 5.
Any other feedback is welcome. Free-form response.
Visualizer
Item Question / Response Options
Instructions This tab visualizes statistics from the full dataset of tutorial dialogues and annotated tutor responses. The dataset contains 300 dialogues with responses from 9 tutors (except for the Novice tutor, who has annotations for 76 dialogues). Participants are asked to explore visualizations for at least 2 tutors.
On a scale from 1 to 5, how informative did you find these visualizations? Scale: 1 = not informative at all, 5 = very informative.
Response options: 1, 2, 3, 4, 5.
On a scale from 1 to 5, how easy was it to use the “Visualizer” tab? Scale: 1 = not easy at all, 5 = very easy.
Response options: 1, 2, 3, 4, 5.
We welcome any further feedback. Free-form response.
C.3 Human Annotation Analysis
We collected 95 annotations for single-tutor responses and 115 for pairwise comparisons. Since annotators were free to choose any dialogues and models, the number of annotations per tutor is not uniform. In pairwise comparisons, Gemini was selected most frequently – 49 out of 115 comparisons included Gemini as one of the tutors. Other tutors appeared in 17-36 comparisons, except for Novice, which was chosen only five times. Overall, Expert responses were preferred most often, winning in 60% of the cases in which the Expert’s response appeared. Sonnet and Llama-3.1-405B were also selected more than half of the time. Novice’s responses never won.
In the single-tutor mode, 37 annotations marked responses as helpful, 31 as not helpful, and 27 as to some extent helpful. The most helpful responses came from Expert, Sonnet, Gemini, and Mistral, each of which was rated helpful in at least half of the corresponding cases. The least helpful responses were from Phi3 and Llama-3.1-8B. Mistral and Llama-3.1-405B were the only tutors without any not helpful annotations, as they had the highest proportion of to some extent helpful ratings.
Appendix D User Interface (UI) Details
Figure 6: Overview of the UI and the Automated Evaluation Tab.
Figure 7: Interface showcasing the selected problem topic with Context, automated Tutor Response, student feedback options, and ground truth verification panel.
Figure 8: UI displaying the enabled Tutor Comparison Mode, allowing users to compare responses from any two selected tutors for the selected problem topic, along with the Context block, selected tutor responses, feedback options, and ground truth verification option.
Figure 9: Displayed Tutor Comparison Visualization Panel showcasing Summary metrics, Spider Chart, Bar Chart, and Differences views for the chosen evaluation results.
Figure 10: Overview of the LLM Evaluation module showcasing the dashboard panel with statistics on topics, conversations, tutors, and evaluation dimensions. The interface also highlights the provided options for Tutor Comparison Mode and Judge Comparison Mode for advanced pedagogical assessment.
Figure 11: User Interface displaying the enabled Tutor Comparison Mode within the LLM Evaluation module, showing the selected problem, tutor responses from two tutors, and the judge LLM for evaluation.
Figure 12: Visualization of the Tutor Performance Summary panel in the Visualizer module, displaying the aggregated evaluation scores for each tutor across all assessment dimensions within the MRBench development dataset.
Figure 13: Interface of the Visualization Controls Panel within the Visualizer module, showing configurable options for selecting tutors and evaluation dimensions to generate comparative spider chart and bar plot visualizations.
Figure 14: Visualization of the Dataset Visualization panel within the Visualizer module, showcasing the spider chart representation of tutor performance across selected evaluation dimensions, based on configurations chosen in the Visualization Controls Panel.
Figure 15: Visualization of the Dataset Visualization panel within the Visualizer module, showcasing the bar plot representation of tutor performance across selected evaluation dimension, based on configurations chosen from the Visualization Controls Panel.